show all calculations work

show all calculations work

d) (450.60+790) / (0.6700)(5070 .)=

06. Convert 0.4890 μm to mile: 

07. Change 0.0062 lb to ng:

08. Transform 8.00 × 10° ft to fl oz: 

09. How many kerneis of popcorn are present in 0.750 pounds of popcorn if the average mass of a kernel of popcorn is 125 mg ? 


10. An unknown object with a mass of 13.75 g raises the volume of water in a graduated cylinder from 25.5 mL to 38.5 mL. What is the density of the unknown object? 

11. The density of methanol, a colorless organic liquid used as solvent is 0.7918 g / mL. Calculate the mass of 128.9 mL of the liquid. 

12. Bromine is a redish-brown liquid. Calculate the volume of 987.55 g of bromine if the density of bromine is 3.12 g/mL. 

13. Ethylene glycol, commonly used as an automobile antifreeze, has a specific gravity of 1.1088 at room temperature (25.0℃). Find the mass of 3.75 L of ethylene glycol at this temperature. 
Answer #1

While solving these questions we will make use of some basic concepts and relationships.

For calculation part we use : BODMAS rule . We will follow the sequence as Bracket, Of, Division, Multiplication, Addition and Subtraction.

1 um = 1 micro meter = 10^-6 meter

1 mile = 1.609 kilometers

1 kilometer = 1000 meter

1 lbs = 453.59 grams

1 gram = 10^9 grams

1 ft^3 = 28.32 liters

1 liter = 33.814 fl oz

For detailed solution and explanation, please refer to images as attached.

PAGE Question a) (450.60 + 790)/(0.6700)( 5070) = ? Solution:- (450.60 +790) (0-6700) (5070) using BODMAS Rule. Solving in SePAGE 3 Question 07 Change 0.0062 16 to ny Solution: - We have О. оо 64 ГБ We know that i | 2b = 453. 54 T and lg = 109 hy whePALE Question (68) Transform 8.00X10-3 ft 3 to fl OZ Here - |ft} t= 28.32 Litece And I litru = 33.814 33.814 fi floz S. 8.00PAGE-6 How Many Kernels of Popcoth are presents in 0.750 pounds, if the Average mass of a Kernel of popcorn is 125 mg? Soluti